The period T of oscillation of a simple pendulum depends on the length L, and acceleration due to gravity g . Determine the exact form of dependence
1
Expert's answer
2020-07-24T10:08:37-0400

The differential equation of oscillation of a simple pendulum


"\frac{d^2\theta}{dt^2}+\frac{g}{l}\sin\theta=0"


after solution


"T=2\pi\sqrt{\frac{l}{g}}\sum{((\frac{(2n)!}{(2^n\cdot n!)})^2\cdot \sin^{2n}\frac{\theta}{2})}."